A diver of mass 65 kg takes off from the 10 m platform from a standing position by jumping upwards with a (near) vertical speed of 2.0 ms-1 at a speed of 30mph.


Sketch graphs of displacement against time and velocity against time for the dive, pointing out the important features of the graphs and the relationship between the two graphs.


1
Expert's answer
2021-03-08T08:23:42-0500

"x_{max}=10+\frac{v^2}{2g}=10.2~m,"

"t_1=\frac vg=0.2~s,"

"t_2=2t_1=0.4~s,"

"t_3=t_1+\sqrt{\frac {2x_ {max} }{g}}=1.6~s,"

"v_{max}=\sqrt{2x_{max}g}=14.3~\frac ms."